OpenStudy (anonymous):

Radicals?? I have no idea how to do this. Rationalize the denominator and simplify. 12/ 7(sqrt) +2

OpenStudy (alekos):

I assume this is 12/(sqrt7 +2) ?

OpenStudy (alekos):

Well?

OpenStudy (anonymous):

12/ (sqrt7) +2 sorry!

OpenStudy (phi):

if you use the equation editor (try it!) , the button in the lower left, you can make it look clearer. do you mean: \[ \frac{12}{\sqrt{7}+2}\] ?

OpenStudy (phi):

if so, you can multiply by the *conjugate* which is a fancy word for multiply top and bottom by \( \sqrt{7}-2\)